### v1

# Exploiting the Structure in Tensor Decompositions for Matrix Multiplication

Abstract.
We present a new algorithm for fast matrix multiplication using tensor decompositions which have special features. Thanks to these features we obtain exponents lower than what the rank of the tensor decomposition suggests. In particular for $6\times 6$ matrix multiplication we reduce the exponent of the recent algorithm by Moosbauer and Poole from $2.8075$ to $2.8016$, while retaining a reasonable leading coefficient.
Bilinear complexity; Strassen’s algorithm; Tensor rank
M. Kauers was supported by the Austrian FWF grants 10.55776/PAT8258123, 10.55776/I6130, and 10.55776/PAT9952223
J. Moosbauer was supported by the Austrian FWF grant 10.55776/PAT8258123
I. Wood was supported by the Austrian FWF grant 10.55776/PAT8258123
conference: ; ; $c$s: Computing methodologies Algebraic algorithms
1. Introduction
Since Strassen’s seminal discovery that two $n\times n$ matrices can be multiplied in $\operatorname{O}(n^{2.81})$ time (STRASSEN, 1969), there has been a race to reduce the upper bound on $\omega$, the exponent of matrix multiplication. Following Strassen’s breakthrough, Bini et al. (Bini et al., 1979) introduced the concept of approximate algorithms and border rank, allowing for a small improvement on the upper bound, finding $\omega\leq 2.78$. This line of research continued to Schönhage’s Asymptotic Sum Inequality (ASI) (Schönhage, 1981), allowing for algorithms that compute disjoint matrix multiplications simultaneously, giving a bound of $\omega\leq 2.52$. Later improvements, based on Strassen’s laser method (Strassen, 1986) and the Coppersmith-Winograd approach (Coppersmith and Winograd, 1990), further reduced the bound on $\omega$ to the currently best upper bound of $\omega<2.371339$ by Alman et al. (Alman et al., 2024).
While there has been a lot of work done to reduce the exponent of matrix multiplication, approximate algorithms (like those used in all the fastest algorithms since 1979) in general either require unreasonable precision or have such large leading coefficients that they are impractical. Most current research into practical algorithms is based on exact algorithms, either trying to reduce the number of multiplications (using, e.g., flip graphs (Kauers and Moosbauer, 2022, 2023a; Kauers and Wood, 2025b; Kauers and Moosbauer, 2023b; Moosbauer and Poole, 2025; Perminov, 2025; Kauers and Wood, 2025a), numerical optimization (Smirnov, 2013; Kaporin, 2024; Novikov et al., 2025a), or reinforcement learning (Fawzi et al., 2022)) or reducing the number of additions (Winograd’s variant of Strassen’s algorithm, published in (Probert, 1976), alternative basis algorithms (Karstadt and Schwartz, 2017; Beniamini and Schwartz, 2019; Schwartz et al., 2024), flip graph search (Perminov, 2025), or common subexpression elimination (Mårtensson and Wagner, 2025)).
Schönhage’s 1981 paper (Schönhage, 1981) uses that an algorithm by Pan (Pan, 1984) has a special property. After 3 iterations, 8 of the recursive calls form a $2\times 2$ matrix multiplication, which can then be computed using Strassen’s algorithm, thereby saving one multiplication. Schwartz and Zwecher (Schwartz and Zwecher, 2025) recently used a similar idea to improve some of Pan’s trilinear aggregation algorithms. The concept also appears in the group theoretic approach to matrix multiplication by Cohn and Umans (Cohn and Umans, 2003; Cohn et al., 2005), who give a variant of the ASI that uses this technique recursively. Schwartz and Zwecher give an explicit algorithm with a $44\times 44$ base case, that achieves the best exponent among exact algorithms with base case smaller than $1000$, while Schönhage’s and Cohn-Umans’ results are purely theoretical bounds on the exponent, with no explicit algorithm provided.
In this paper we propose a new recursive matrix multiplication algorithm that uses the same idea. If some recursive calls share one of the inputs or have an output that is used in multiple positions, then they are treated as a single matrix multiplication of larger size. Applying this technique repeatedly improves the exponent of a matrix multiplication algorithm without reducing the number of multiplications over the base field. In particular, we improve the recent $6\times 6$ matrix multiplication algorithm by Moosbauer and Poole (Moosbauer and Poole, 2025) from an exponent of $2.8075$ to $2.8016$, compared to $2.8073$ of Strassen’s algorithm. Our algorithm outperforms the standard algorithm for $n\geq 1000$ in terms of total operation count. This does not necessarily mean that in an implementation one would already observe a speedup for such matrix sizes, but it clearly puts it in the realm of practical algorithms by Pan’s definition (Pan, 1984), who generously drew the line at $n=10^{20}$.
We ran a search for such algorithms, first using a flip graph search to find algorithms using a minimal number of multiplications while retaining the special structure needed to apply our new recursive method. Then we optimize the number of additions using DeGroote actions. These steps produce an algorithm that is valid only modulo $2$, thus we apply Hensel lifting to lift the algorithm to work over the integers, and finally we reduce the number of additions further by Mårtensson and Wagner’s common subexpression elimination method (Mårtensson and Wagner, 2025). We find improvements in the exponent for several small base cases including $3\times 3$ matrix multiplication, where we reduce the exponent from Laderman’s $2.854$ (Laderman, 1976) to $2.836$.
2. Background
Let $K$ be a field and $R$ be a $K$-algebra, and let $\mathbf{A}\in R^{n\times m},\mathbf{B}\in R^{m\times p}$.
Strassen’s algorithm (STRASSEN, 1969) uses a divide and conquer strategy to multiply matrices faster than the standard algorithm. First, it reduces the multiplication of two $n\times n$ matrices to multiplying two $2\times 2$ matrices whose entries are $\frac{n}{2}\times\frac{n}{2}$ matrices using block matrix multiplication. Then it computes this $2\times 2$ block matrix product using only $7$ multiplications of the subblocks by linearly combining them in a clever way. Using this strategy recursively gives an algorithm to multiply $n\times n$ matrices using $\operatorname{O}(n^{\log_{2}(7)})$ operations in $R$. The exponent only depends on the number of multiplications, since the cost of the recursive call dominates the cost of the linear combinations.
Usually, these algorithms are written using the language of tensors.
Definition 1.
Let $n,m,p\in\mathbb{N}$ and $V_{1}=R^{nm},V_{2}=R^{mp},V_{3}=R^{pn}$. Denote by $\{a_{ij}:1\leq i\leq n,1\leq j\leq m\}$ the standard basis of $V_{1}$, $\{b_{jk}:1\leq j\leq m,1\leq k\leq p\}$ the standard basis of $V_{2}$ and $\{c_{ki}:1\leq k\leq p,1\leq i\leq n\}$ the standard basis of $V_{3}$. We define the matrix multiplication tensor to be
$\langle n,m,p\rangle=\sum_{i,j,k=1}^{n,m,p}a_{ij}\otimes b_{jk}\otimes c_{ki}\in V_{1}\otimes V_{2}\otimes V_{3}.$
This tensor encodes matrix multiplication in the following sense: let $A,B,C$ be matrix spaces over $R$ with dimensions $n\times m,m\times p,n\times p$ respectively and denote by $\operatorname{Bil}(A,B;C)$ the space of all bilinear maps from $A\times B$ to $C$. Then we can define an isomorphism $\Phi\colon V_{1}\otimes V_{2}\otimes V_{3}\to\operatorname{Bil}(A,B;C)$ by
$\Phi(a_{i_{2}j_{1}}\otimes b_{j_{2}k_{1}}\otimes c_{k_{2}i_{1}})(\mathbf{A},\mathbf{B})=\mathbf{A}_{i_{2}j_{1}}\mathbf{B}_{j_{2}k_{1}}\varGamma_{i_{1}k_{2}}$
where $\mathbf{A}\in A,\mathbf{B}\in B$, $\mathbf{A}_{ij},\mathbf{B}_{ij}$ represent the $(i,j)$-th entry of $\mathbf{A}$ and $\mathbf{B}$ respectively and $\varGamma_{ij}\in C$ is the matrix with all zeros except for a $1$ at the $(i,j)$-th entry. Note the $c_{ij}$ corresponds to $\mathbf{C}_{ji}$ in the product matrix; this is a standard convention to keep the cyclic nature of matrix multiplication.
Definition 2.
We call a tensor $T\in U_{1}\otimes U_{2}\otimes U_{3}$ a restriction of $T^{\prime}\in U_{1}^{\prime}\otimes U_{2}^{\prime}\otimes U_{3}^{\prime}$, denoted by $T\leq T^{\prime}$, if there are homomorphisms $\phi_{1}\colon U_{1}^{\prime}\to U_{1},\phi_{2}\colon U_{2}^{\prime}\to U_{2},\phi_{3}\colon U_{3}^{\prime}\to U_{3}$ with $T=(\phi_{1}\otimes\phi_{2}\otimes\phi_{3})T^{\prime}$. If we need to refer to the specific homomorphism $\phi=(\phi_{1}\otimes\phi_{2}\otimes\phi_{3})$, we write $T\leq_{\phi}T^{\prime}$.
If there are such isomorphisms $\phi_{1},\phi_{2},\phi_{3}$, we call $T$ and $T^{\prime}$ isomorphic and write $T\cong T^{\prime}$.
We write the unit tensor of size $k$ as $\langle k\rangle=\sum_{i=1}^{k}e_{i}\otimes e_{i}\otimes e_{i}\in R^{k}\otimes R^{k}\otimes R^{k}$ where $\{e_{1},\dots,e_{k}\}$ is the standard basis for $R^{k}$.
If a tensor $T$ is a restriction of a tensor $T^{\prime}$ this means that any algorithm to compute $\Phi(T^{\prime})$ can be transformed into one to compute $\Phi(T)$ by applying the linear maps $\phi_{1},\phi_{2},\phi_{3}$ to the inputs and output. In particular, if we have a restriction of the form $\langle n,m,p\rangle\leq\langle r\rangle$, then we have an algorithm to multiply $n\times m$ and $m\times p$ matrices using $r$ multiplications in $R$. A restriction of the form $T\leq\langle r\rangle$ gives rise to a decomposition of $T$ into $r$ summands of the form
$T=\sum_{i=1}^{r}a_{i}\otimes b_{i}\otimes c_{i}.$
We then call $r$ the rank of this decomposition and a tensor that can be written in the form $a\otimes b\otimes c$ a rank-one tensor. The rank of a tensor $T$ is defined as the smallest integer $r$ such that $T$ has a decomposition of rank $r$.
Together with Strassen’s recursive block matrix multiplication strategy, this gives the following fundamental result.
Theorem 3.
If $\langle n,m,p\rangle\leq\langle r\rangle$ then there is an algorithm to multiply $N\times N$ matrices in $\operatorname{O}(N^{3\log_{nmp}(r)})$ operations in $R$.
Using the decomposition Strassen found, we get $\langle 2,2,2\rangle\leq\langle 7\rangle$ and hence $\omega\leq\log_{2}(7)$. Further analysis allows to compute the leading coefficient of this algorithm as well, as we will see next. While the exponent is given by the rank of a decomposition, i.e. the number of multiplications, the leading coefficient depends on the number of additions and scalar multiplications needed in the restriction. We will denote this number by $A(\phi)$ if the restriction is given by $\phi$. The homomorphism for a restriction is in general not unique, but in our case every restriction is witnessed by an explicit homomorphism, so we just write $A$ when the homomorphism is clear from the context.
Using this definition, Strassen’s original algorithm has $A=18$. Using this we can compute the leading coefficient of the algorithm, which is done in general in Theorem 4. Although this result is well-known, we present its proof in detail since as it establishes the structure for the more complex proof in Section 3.
Theorem 4.
For a restriction $\langle n,n,n\rangle\leq_{\phi}\langle r\rangle$ with $\omega_{0}=\log_{n}(r)<3$, the algorithm from Theorem 3 takes at most
$\left(2(n-1)^{3-\omega_{0}}+\frac{r(2^{\omega_{0}}-1)+4A(\phi)}{r-n^{2}}(n-1)^{2-\omega_{0}}\right)N^{\omega_{0}}+\operatorname{O}(N^{2})$
operations to multiply $N\times N$ matrices.
Proof.
We denote by $T(N)$ the total number of operations the algorithm needs to multiply $N\times N$ matrices. For $1\leq N<n$ we use the standard algorithm, so $T(N)=2N^{3}-N^{2}$. For $N\geq n$, we compute $r$ recursive multiplications of size $\lceil N/n\rceil\times\lceil N/n\rceil$, and need $A\lceil N/n\rceil^{2}$ operations to form the linear combinations. The ceiling accounts for zero padding to the next multiple of $n$. Hence, we have the recurrence relation
$T(N)=rT(\lceil N/n\rceil)+A\lceil N/n\rceil^{2}$
for all $N\geq n$.
We will show by induction on $N$ that
$T(N)\leq LN^{\omega_{0}}-dN^{2}$
for $L=2(n-1)^{3-\omega_{0}}+\frac{r(2^{\omega_{0}}-1)+4A}{r-n^{2}}(n-1)^{2-\omega_{0}}$ and $d$ to be determined later. Suppose the statement holds for all $N<k$. Then we have
$T(k)$ $=rT(\lceil k/n\rceil)+A\lceil k/n\rceil^{2}$
$\leq r(L\lceil k/n\rceil^{\omega_{0}}-d\lceil k/n\rceil^{2})+A\lceil k/n\rceil^{2}$
$\leq r(L(k/n+1)^{\omega_{0}}-d(k/n)^{2})+A(k/n+1)^{2}$
$\leq r(L(k/n)^{\omega_{0}}+c(k/n)^{\omega_{0}-1}-d(k/n)^{2})+A(k/n+1)^{2}$
$\leq Lk^{\omega_{0}}+\frac{rc}{n^{2}}k^{2}-\frac{rd}{n^{2}}k^{2}+\frac{4A}{n^{2}}k^{2},$
where $c=2^{\omega_{0}}-1$ is a chosen such that $(x/n+1)^{\omega_{0}}\leq(x/n)^{\omega_{0}}+c(x/n)^{\omega_{0}-1}$ for all $x\geq n$. Choosing $d=\frac{rc+4A}{r-n^{2}}$ completes the induction step.
It remains to show that for all $N<n$ we have $T(N)\leq LN^{\omega_{0}}-dN^{2}$. Since $T(N)=2N^{3}-N^{2}$ for $N<n$, it suffices to show that $L\geq 2N^{3-\omega_{0}}+(d-1)N^{2-\omega_{0}}$ for all $1\leq N<n$, which is the case. ∎
In the case of Strassen’s original algorithm, which used a restriction $\langle 2,2,2\rangle\leq_{\phi}\langle 7\rangle$ with $A(\phi)=18$, we find that this theorem gives us a bound on the leading coefficient $L\leq 40$. Usually the leading coefficient bounds are computed in the idealized case where every recursive call is assumed to be done exactly. In this case, the leading coefficient of Strassen’s algorithm is $7$, the general formula being $L\leq\frac{A}{r-n^{2}}+1$ (Karstadt and Schwartz, 2017).
Theorem 3 is very important, however, all improvements since 1981 to the upper bound on $\omega$ have come from the following stronger theorem due to Schönhage.
Theorem 5 (Asymptotic Sum Inequality (ASI)).
If
$\bigoplus_{i=1}^{k}\langle n_{i},m_{i},p_{i}\rangle\leq\langle r\rangle,$
and $\omega_{0}$ is such that
$\sum_{i=1}^{k}(n_{i}m_{i}p_{i})^{\frac{\omega_{0}}{3}}=r,$
then for every $\epsilon>0$, there exists an algorithm to multiply $n\times n$ matrices using $\operatorname{O}(n^{\omega_{0}+\epsilon})$ operations in $R$.
Our result is related to the following generalization of the ASI, which was first stated by Cohn and Umans (Cohn and Umans, 2003).
Corollary 6.
If
$\bigoplus_{i=1}^{k}\langle n_{i},m_{i},p_{i}\rangle\leq\bigoplus_{i=k+1}^{k+q}\langle n_{i},m_{i},p_{i}\rangle,$
and $\omega_{0}$ is such that
$\sum_{i=1}^{k}(n_{i}m_{i}p_{i})^{\frac{\omega_{0}}{3}}=\sum_{i=k+1}^{k+q}(n_{i}m_{i}p_{i})^{\frac{\omega_{0}}{3}},$
then for every $\epsilon>0$, there exists an algorithm to multiply $n\times n$ matrices using $\operatorname{O}(n^{\omega_{0}+\epsilon})$ operations in $R$.
While the ASI has proved very useful for finding ever tighter bounds on the exponent of matrix multiplication, it has, to our knowledge, never been used for finding fast matrix multiplication algorithms outside the purely asymptotic regime. The theorem is frequently used in the context of approximate algorithms, requiring either an unreasonable level of precision for most applications or an extremely large leading coefficient, as well as very large base cases, rendering them functionally useless in the context of practical algorithms.
In this work we do not aim to improve the exponent of matrix multiplication in general, but to find algorithms which could be used in practical computations, while achieving a better complexity. This does not mean that the algorithms presented here are practical in the sense of being faster than currently used algorithms for matrix sizes that appear in practical computations, but at least they could reasonably be implemented and used for matrix sizes that fit into the memory of a large computer.
We will use direct sums and Kronecker products of tensors in the following part. The important property is that they are compatible with the notion of restriction, so we have
$T_{1}\leq T_{1}^{\prime},T_{2}\leq T_{2}^{\prime}\implies T_{1}\oplus T_{2}\leq T_{1}^{\prime}\oplus T_{2}^{\prime},T_{1}\otimes T_{2}\leq T_{1}^{\prime}\otimes T_{2}^{\prime}.$
Readers less familiar with these concepts can view these operations as follows: The direct sum $\oplus$ corresponds to computing two matrix multiplications independently, while the Kronecker product $\otimes$ corresponds to the inner and outer part of a block matrix multiplication. In particular, we have that
$\langle n_{1},m_{1},p_{1}\rangle\otimes\langle n_{2},m_{2},p_{2}\rangle\cong\langle n_{1}n_{2},m_{1}m_{2},p_{1}p_{2}\rangle.$
Since we will frequently encounter direct sums and Kronecker products of matrix multiplication tensors of the same format, we use the notation
$k\odot\langle n,m,p\rangle=\bigoplus_{i=1}^{k}\langle n,m,p\rangle,\quad\langle n,m,p\rangle^{\otimes k}=\bigotimes_{i=1}^{k}\langle n,m,p\rangle.$
We will also use the well-known fact that one can permute the dimensions of matrix multiplication tensors. Specifically, we get from $(\mathbf{A}\mathbf{B})^{T}=\mathbf{A}^{T}\mathbf{B}^{T}$ that we transform any algorithm from a restriction of $\langle n,m,p\rangle$ into one for $\langle p,m,n\rangle$ and from the cyclic symmetry of the matrix multiplication tensor it follows that we can transform it to $\langle m,p,n\rangle$ as well.
3. divide less, conquer more
While Schönhage’s ASI is an existence result, we present an explicit algorithm with the asymptotic complexity given by the generalized ASI. In contrast to the classical ASI we do not allow a direct sum of matrix multiplication tensors on the left side, but only on the right side. This means that we decompose a matrix multiplication tensor into a sum of smaller matrix multiplications, not just rank-one tensors. This allows to divide a matrix multiplication into larger subblocks, thereby performing more recursive steps (and hence more savings compared to the standard algorithm).
The concept of achieving additional savings this way originates from Schönhage’s 1981 paper (Schönhage, 1981), in which it is shown to find improvements to Pan’s use of the ASI (Pan, 1984) by taking tensor powers and then applying Strassen’s algorithm. We will illustrate the idea using the decomposition of $\langle 6,6,6\rangle$ found by Moosbauer and Poole (Moosbauer and Poole, 2025). They show that $\langle 6,6,6\rangle$ can be decomposed into a sum of $153$ rank-one tensors, analyzing the structure of this decomposition, we can find the restriction $\langle 6,6,6\rangle\leq 137\odot\langle 1\rangle\oplus 8\odot\langle 1,1,2\rangle$. Applying a cyclic permuation $\sigma$ to this restriction gives $\langle 6,6,6\rangle\leq 137\odot\langle 1\rangle\oplus 8\odot\langle 2,1,1\rangle$, and $\langle 6,6,6\rangle\leq 137\odot\langle 1\rangle\oplus 8\odot\langle 1,2,1\rangle$. If we then consider the Kronecker product of these three restrictions we get
$\langle 6,6,6\rangle^{\otimes 3}$ $\leq 137^{3}\odot\langle 1\rangle\oplus 8\cdot 137^{2}\odot(\langle 1,1,2\rangle\oplus\langle 2,1,1\rangle\oplus\langle 1,2,1\rangle)$
$\kern-1.0pt\oplus 8^{2}\cdot 137\odot(\langle 1,2,2\rangle\oplus\langle 2,1,2\rangle\oplus\langle 2,2,1\rangle)\oplus 8^{3}\odot\langle 2,2,2\rangle.$
While we have to compute $\langle 1,1,2\rangle,\langle 1,2,2\rangle$ and their cyclic permutations by the standard algorithm, using $2$ and $4$ multiplications respectively, we can compute $\langle 2,2,2\rangle$ using Strassen’s algorithm with $7$ multiplications. In the language of tensor restrictions we use $\langle 1,1,2\rangle\leq\langle 2\rangle$, $\langle 1,2,2\rangle\leq\langle 4\rangle$ and $\langle 2,2,2\rangle\leq\langle 7\rangle$, to conclude that $\langle 216,216,216\rangle\leq\langle 3581065\rangle$. Applying Theorem 3 to this restriction gives an algorithm with exponent $\omega_{0}=2.80751$, which is only slightly smaller than the original exponent $2.80754$. We can do even better by taking higher tensor powers.
But this is not all. The key observation is that instead of using Strassen’s decomposition to save additional multiplications, we can use any fast matrix multiplication algorithm for this recursive calls, in particular we can use the same strategy recursively. In this example we would get an algorithm with exponent $\omega_{0}=2.80496$.
We will show that we can use this process in a recursive matrix multiplication algorithm that is very similar to the standard Strassen-like algorithms obtained from tensor decompositions that has the complexity as given by the generalized ASI while retaining a reasonable leading coefficient.
In Algorithm 1 we write $\mathcal{L}_{1}^{(i,j)},\mathcal{L}_{2}^{(i,j)},\mathcal{L}_{3}^{(k,l)}$ for the functions that compute the linear combinations of subblocks as given by the tensor restriction. Unlike in Strassen’s algorithm, some of these block matrices are recombined into larger matrices of size $n_{i}\times m_{i}$ and $m_{i}\times p_{i}$ respectively.
Algorithm 1 Matrix Mult for a given restriction $\langle n,m,p\rangle\leq\bigoplus_{i=0}^{q-1}s_{i}\odot\langle n_{i},m_{i},p_{i}\rangle$
1:An $N\times M$ matrix $\mathbf{A}$ and an $M\times P$ matrix $\mathbf{B}$, a treshold $N_{0}$.
2:The matrix product $\mathbf{C}=\mathbf{A}\mathbf{B}$
3:function MatrixMults($\mathbf{A},\mathbf{B},N,M,P$)
4:  if $\min(N,M,P)\leq N_{0}$ then
5:   return $\mathbf{A}\mathbf{B}$ computed through standard algorithm
6:  end if
7:  if $n\nmid N\lor m\nmid M\lor p\nmid P$ then
8:   $N\leftarrow\lceil N/n\rceil n$
9:   $M\leftarrow\lceil M/m\rceil m$
10:   $P\leftarrow\lceil P/p\rceil p$
11:   pad $\mathbf{A}$ and $\mathbf{B}$ with zeros to dimensions $N\times M$ and $M\times P$
12:  end if
13:  for $i\leftarrow 0$ to q1q-1 do
14:   for $j\leftarrow 0$ to $s_{i}-1$ do
15:     $X\leftarrow\mathcal{L}_{1}^{(i,j)}(\mathbf{A})$
16:     $Y\leftarrow\mathcal{L}_{2}^{(i,j)}(\mathbf{B})$
17:     $Z_{i,j}\leftarrow\textsc{MatrixMult}(X,Y,Nn_{i}/n,Nm_{i}/m,Pp_{i}/p)$
18:   end for
19:  end for
20:  for $k\leftarrow 0$ to N/n1N/n-1 do
21:   for $l\leftarrow 0$ to P/p1P/p-1 do
22:     $\mathbf{C}_{k,l}\leftarrow\mathcal{L}_{3}^{(k,l)}(\mathbf{Z})$
23:   end for
24:  end for
25:  return $\mathbf{C}$
26:end function
Algorithm 1 correctly computes the matrix product $\mathbf{C}$, since it computes all the same matrix products as the standard Strassen-like algorithm obtained from the tensor decomposition. However, our algorithm computes some of them simultaneously as a larger matrix.
If in Algorithm 1 some of the recursive calls are rectangular matrix multiplications (as will be the case in all our results), then for some recursive paths the number of recursions is limited to $\min(\log_{\frac{n_{i}}{n}}(N),\log_{\frac{m_{i}}{m}}(M),\log_{\frac{p_{i}}{p}}(P))$, before we encounter a subproblem of shape $\langle N^{\prime},M^{\prime},1\rangle$, or a permutation thereof, which forces us to switch to the standard algorithm. However, we do not need to use the same restriction in each recursive call. Similar to the example discussed above we can alternate between three cyclic permutations of a restriction, to ensure that most recursive paths can continue for longer.
4. Complexity Analysis
We will now analyze the complexity of Algorithm 1, showing that it achieves the complexity promised by the generalized ASI. We start by analyzing the idealized case where all recursive calls are square matrix multiplications.
Theorem 1.
If $N=M=P$, $n=m=p$ and $n_{i}=m_{i}=p_{i}$ for all $0\leq i<q$, then there exists an $N_{0}$ such that Algorithm 1 takes $\operatorname{O}(N^{\omega_{0}})$ operations to compute the product of $N\times N$ matrices, where $\omega_{0}$ is such that $n^{\omega_{0}}=\sum_{i=0}^{q-1}s_{i}n_{i}^{\omega_{0}}$.
The proof of this theorem will follow very closely the proof of Theorem 4.
Proof.
Let $T(N)$ be the number of operations Algorithm 1 needs to multiply $N\times N$ matrices.
We choose $n^{\prime}\in\mathbb{N}$ such that $n^{\prime}>c=\frac{(2^{\omega_{0}}-1)n}{\sum_{i=0}^{q-1}s_{i}\left(\frac{n_{i}}{n}\right)^{\omega_{0}-1}-1}$ and we choose $N_{0}=nn^{\prime}-1$.
Then we have that $T(N)=2N^{3}-N^{2}$ for all $1\leq N<nn^{\prime}$ and
$T(N)=\sum_{i=0}^{q-1}s_{i}T(\lceil N/n\rceil n_{i})+A\lceil N/n\rceil^{2},$
for all $N\geq nn^{\prime}$.
Let
$d=\frac{A/n^{2}}{\sum_{i=0}^{q-1}s_{i}\left(\frac{n_{i}}{n}\right)^{2}-1}$
and
$L=\max_{n^{\prime}\leq x<nn^{\prime}}\frac{2x^{4-\omega_{0}}+(d-1)x^{3-\omega_{0}}}{x-c}.$
Since $n^{\prime}>c$, we can ensure that the denominator is positive, so $L$ is always well defined.
We will prove by induction on $N$ that
(1) $T(N)\leq LN^{\omega_{0}}-LcN^{\omega_{0}-1}-dN^{2}$
for all $N\geq n^{\prime}$.
For the base case we have that
$T(N)=2N^{3}-N^{2}\leq LN^{\omega_{0}}-LcN^{\omega_{0}-1}-dN^{2}$
for all $n^{\prime}\leq N<nn^{\prime}$ by our choice of $L$.
For the induction step let $k\geq nn^{\prime}$ be such that (1) holds for all $n^{\prime}\leq N<k$. Let $M=\lceil\frac{N}{n}\rceil$, then we have the inequalities $\frac{N}{n}\leq M\leq\frac{N}{n}+1$ and $M\leq 2\frac{N}{n}$. We will also use that $(x+1)^{\omega_{0}}\leq x^{\omega_{0}}+(2^{\omega_{0}}-1)x^{\omega_{0}-1}$ for all $x\geq 1$. Then we have
$T(N)$ ${}\leq{}$ $\sum_{i=0}^{q-1}s_{i}T(\lceil\frac{N}{n}\rceil n_{i})+A\lceil\frac{N}{n}\rceil^{2}$
$\leq{}$ $\sum_{i=0}^{q-1}s_{i}\left(L(Mn_{i})^{\omega_{0}}-Lc(Mn_{i})^{\omega_{0}-1}-d(Mn_{i})^{2}\right)+AM^{2}$
$\leq{}$ $L(Mn)^{\omega_{0}}-LcM^{\omega_{0}-1}\sum_{i=0}^{q-1}s_{i}n_{i}^{\omega_{0}-1}-M^{2}\left(d\sum_{i=0}^{q-1}s_{i}n_{i}^{2}-A\right)$
$\leq{}$ $L\left(\frac{N}{n}+1\right)^{\omega_{0}}n^{\omega_{0}}-Lc\left(\frac{N}{n}\right)^{\omega_{0}-1}\sum_{i=0}^{q-1}s_{i}n_{i}^{\omega_{0}-1}$
$-\left(\frac{N}{n}\right)^{2}\left(d\sum_{i=0}^{q-1}s_{i}n_{i}^{2}-A\right)$
$\leq{}$ $LN^{\omega_{0}}-L\left(-(2^{\omega_{0}}-1)n+c\sum_{i=0}^{q-1}s_{i}\left(\frac{n_{i}}{n}\right)^{\omega_{0}-1}\right)N^{\omega_{0}-1}$
$-\left(d\sum_{i=0}^{q-1}s_{i}\left(\frac{n_{i}}{n}\right)^{2}-\frac{A}{n^{2}}\right)N^{2}$
$\leq{}$ $LN^{\omega_{0}}-LcN^{\omega_{0}-1}-dN^{2},$
where the last step follows from our choice of $c$ and $d$. ∎
As was the case in Theorem 3, the leading coefficient $L$ given here is a strict upper bound. However, to make the leading coefficient more comparable to those usually presented for standard Strassen-like algorithms, we recommend disregarding the cost of zero padding in our algorithm. This gives a different coefficient for the algorithm, a leading coefficient which is likely never attained, but is close enough to serve as a point of comparison. This coefficient is given by
(2) $L=\frac{A}{-n^{2}+\sum_{i=0}^{q-1}s_{i}n_{i}^{2}}+1,$
and when applied to Strassen’s algorithm ($n=2,q=1,n_{0}=1,s_{0}=7,A=18$) this gives $L=7$ as expected.
As pointed out at the end of Section 3, to achieve the claimed complexity for more general decompositions we need to alternate between different decompositions. We will first argue that taking Kronecker products and Kronecker powers of a given restriction and its permutations and then applying Algorithm 1 is essentially equivalent to using Algorithm 1 with appropriate permutations of the original restriction.
Indeed we have that one iteration of Algorithm 1 using the $k$-th tensor power of a given restriction performs essentially the same computation as taking exactly $k$ recursive steps of the algorithm when using the original restriction. The only difference is that for the tensor power we need to do all zero padding at the beginning, while for the original restriction we do zero padding at each recursive step. The total size of the zero padding is the same in both cases, but doing it in every iteration gives a lower operation count for computing $\mathcal{L}_{1},\mathcal{L}_{2}$ and $\mathcal{L}_{3}$, so this does not affect the exponent.
In our analysis we will also assume that we have a restriction of the form
$\langle n,n,n\rangle\leq\bigoplus_{i=0}^{q_{0}-1}s_{i}\odot(\langle n_{i},m_{i},p_{i}\rangle\oplus\langle m_{i},p_{i},n_{i}\rangle\oplus\langle p_{i},n_{i},m_{i}\rangle).$
To bring a given restriction into this form we can first take a tensor product of three cyclic permutations of a restriction. This ensures that for all $i$ where not $n_{i}=m_{i}=p_{i}$ we have that all cyclic permutations appear with the same coefficient. To deal with the cases where $n_{i}=m_{i}=p_{i}$ we write $\langle n_{i},n_{i},n_{i}\rangle=\frac{1}{3}\odot(\langle n_{i},n_{i},n_{i}\rangle\oplus\langle n_{i},n_{i},n_{i}\rangle\oplus\langle n_{i},n_{i},n_{i}\rangle)$, note that this is only notation, as $\odot$ is not defined for non-integer factors, but all the proofs go through unchanged.
Theorem 2.
Suppose
$\langle n,n,n\rangle\leq\bigoplus_{i=0}^{q-1}s_{i}\odot(\langle n_{i},m_{i},p_{i}\rangle\oplus\langle m_{i},p_{i},n_{i}\rangle\oplus\langle p_{i},n_{i},m_{i}\rangle).$
Let $\omega_{k}$ be the exponent achieved by applying Algorithm 1 to the $k$-th tensor power of this decomposition. Then the $\omega_{k}$ form a decreasing sequence with
$\lim_{k\to\infty}\omega_{k}=\omega_{0},$
where $\omega_{0}$ is such that
$n^{\omega_{0}}=3\sum_{i=0}^{q-1}s_{i}(n_{i}m_{i}p_{i})^{\omega_{0}/3}.$
Proof.
Consider the $k$-th tensor power of the initial restriction. Since the tensor product distributes over the direct sum, we obtain
$\langle n^{k},n^{k},n^{k}\rangle\cong\langle n,n,n\rangle^{\otimes k}$
$\qquad\leq\left(\bigoplus_{i=0}^{q-1}s_{i}\odot(\langle n_{i},m_{i},p_{i}\rangle\oplus\langle m_{i},p_{i},n_{i}\rangle\oplus\langle p_{i},n_{i},m_{i}\rangle)\right)^{\otimes k}$
$\qquad=\bigoplus_{r_{0}+r_{1}+\dots+r_{3k-1}=k}S_{\mathbf{r}}\odot\langle N_{\mathbf{r}},M_{\mathbf{r}},P_{\mathbf{r}}\rangle$
where
$S_{\mathbf{r}}$ $=\binom{k}{r_{0},\dots,r_{3k-1}}\prod_{i=0}^{q-1}s_{i}^{r_{3i}+r_{3i+1}+r_{3i+2}},$ $\quad N_{\mathbf{r}}$ $=\prod_{i=0}^{q-1}n_{i}^{r_{3i}}m_{i}^{r_{3i+1}}p_{i}^{r_{3i+2}}$
$M_{\mathbf{r}}$ $=\prod_{i=0}^{q-1}m_{i}^{r_{3i}}p_{i}^{r_{3i+1}}n_{i}^{r_{3i+2}},$ $P_{\mathbf{r}}$ $=\prod_{i=0}^{q-1}p_{i}^{r_{3i}}n_{i}^{r_{3i+1}}m_{i}^{r_{3i+2}}.$
Since Theorem 1 requires a decomposition into square schemes, we now restrict each $\langle N_{\mathbf{r}},M_{\mathbf{r}},P_{\mathbf{r}}\rangle$ to a direct sum of square matrix multiplication tensors. Let $D_{\mathbf{r}}=\min(N_{\mathbf{r}},M_{\mathbf{r}},P_{\mathbf{r}})$. We have that
$\langle N_{\mathbf{r}},M_{\mathbf{r}},P_{\mathbf{r}}\rangle\leq{}$ $\lfloor\frac{N_{\mathbf{r}}}{D_{\mathbf{r}}}\rfloor\lfloor\frac{M_{\mathbf{r}}}{D_{\mathbf{r}}}\rfloor\lfloor\frac{P_{\mathbf{r}}}{D_{\mathbf{r}}}\rfloor\odot\langle D_{\mathbf{r}},D_{\mathbf{r}},D_{\mathbf{r}}\rangle$
$\oplus(N_{\mathbf{r}}M_{\mathbf{r}}P_{\mathbf{r}}-\lfloor\frac{N_{\mathbf{r}}}{D_{\mathbf{r}}}\rfloor\lfloor\frac{M_{\mathbf{r}}}{D_{\mathbf{r}}}\rfloor\lfloor\frac{P_{\mathbf{r}}}{D_{\mathbf{r}}}\rfloor D_{\mathbf{r}}^{3})\odot\langle 1\rangle.$
If we apply Algorithm 1 using this restriction, then Theorem 1 gives
$n^{k\omega_{k}}\leq{}$ $\sum_{\mathbf{r}}S_{\mathbf{r}}\lfloor\frac{N_{\mathbf{r}}}{D_{\mathbf{r}}}\rfloor\lfloor\frac{M_{\mathbf{r}}}{D_{\mathbf{r}}}\rfloor\lfloor\frac{P_{\mathbf{r}}}{D_{\mathbf{r}}}\rfloor D_{\mathbf{r}}^{\omega_{k}}$
${}+S_{\mathbf{r}}\left(N_{\mathbf{r}}M_{\mathbf{r}}P_{\mathbf{r}}-\lfloor\frac{N_{\mathbf{r}}}{D_{\mathbf{r}}}\rfloor\lfloor\frac{M_{\mathbf{r}}}{D_{\mathbf{r}}}\rfloor\lfloor\frac{P_{\mathbf{r}}}{D_{\mathbf{r}}}\rfloor D_{\mathbf{r}}^{3}\right)$
$\leq{}$ $\sum_{\mathbf{r}}S_{\mathbf{r}}N_{\mathbf{r}}M_{\mathbf{r}}P_{\mathbf{r}}D_{\mathbf{r}}^{\omega_{0}-3}+S_{\mathbf{r}}N_{\mathbf{r}}M_{\mathbf{r}}P_{\mathbf{r}}$
$\leq{}$ $\sum_{\mathbf{r}}S_{\mathbf{r}}N_{\mathbf{r}}M_{\mathbf{r}}P_{\mathbf{r}}(D_{\mathbf{r}}^{\omega_{k}-3}+1)$
$\leq{}$ $2\sum_{\mathbf{r}}S_{\mathbf{r}}N_{\mathbf{r}}M_{\mathbf{r}}P_{\mathbf{r}}D_{\mathbf{r}}^{\omega_{k}-3}$
$\leq{}$ $2\sum_{\mathbf{r}}S_{\mathbf{r}}N_{\mathbf{r}}M_{\mathbf{r}}P_{\mathbf{r}}(N_{\mathbf{r}}M_{\mathbf{r}}P_{\mathbf{r}})^{\frac{\omega_{k}-3}{3}}$
$={}$ $2\sum_{\mathbf{r}}S_{\mathbf{r}}(N_{\mathbf{r}}M_{\mathbf{r}}P_{\mathbf{r}})^{\frac{\omega_{k}}{3}}$
$={}$ $2\left(3\sum_{i=0}^{q-1}s_{i}(n_{i}m_{i}p_{i})^{\frac{\omega_{k}}{3}}\right)^{k}$
Taking the $k$-th root, we get $n^{\omega_{k}}\leq 2^{\frac{1}{k}}\sum_{i=0}^{q-1}s_{i}(n_{i}m_{i}p_{i})^{\frac{\omega_{k}}{3}}$. Taking $k\rightarrow\infty$ gives the claimed result, since $2^{\frac{1}{k}}\rightarrow 1$. ∎
Our final bound actually gives us a concrete idea of how quickly these $\omega_{k}$ approach $\omega_{0}$. Consider the function
$f(w)=\log\left(3\sum_{i=0}^{q-1}s_{i}(n_{i}m_{i}p_{i})^{\frac{w}{3}}\right)-w\log n.$
We have $f(\omega_{0})=0$, and by the mean value theorem, for some $\xi$ between $\omega_{0}$ and $\omega_{k}$ we have
$f(\omega_{k})=f(\omega_{0})+f^{\prime}(\xi)(\omega_{k}-\omega_{0})=f^{\prime}(\xi)(\omega_{k}-\omega_{0}).$
We find
$f^{\prime}(w)=\frac{\sum_{i=0}^{q-1}s_{i}(n_{i}m_{i}p_{i})^{\frac{w}{3}}\log\frac{n_{i}m_{i}p_{i}}{n^{3}}}{3\sum_{i=0}^{q-1}s_{i}(n_{i}m_{i}p_{i})^{\frac{w}{3}}}$
and note that this must be negative since we have $n_{i}m_{i}p_{i}<n^{3}$ for all $0\leq i<q$ in any restriction that gives a non-trivial bound on $\omega_{0}$. From the proof of Theorem 2 follows that $f(\omega_{k})\geq\frac{-\log 2}{k}$. Thus, we have that
$\frac{-\log 2}{k}\leq f(\omega_{k})=f^{\prime}(\xi)(\omega_{k}-\omega_{0})$
so
$\omega_{0}+\frac{-\log 2}{kf^{\prime}(\xi)}\geq\omega_{k}.$
Let $F=\inf_{x\in(\omega_{0},\omega_{1})}|f^{\prime}(x)|$, then we have $\omega_{k}\leq\omega_{0}+\frac{\log 2}{Fk}$ for all $k\geq 1$.
The bound on the leading coefficient of this algorithm depends on $k$. To quantify this dependence we analyze the number of additions and scalar multiplications performed by such an algorithm.
Lemma 3.
If $\phi^{\otimes k}$ is the homomorphism witnessing the $k$-th tensor power of a restriction
$\langle n,n,n\rangle\leq_{\phi}\bigoplus_{i=0}^{q-1}s_{i}\odot\langle n_{i},m_{i},p_{i}\rangle,$
then we have
$A(\phi^{\otimes k})\leq A(\phi)\frac{r^{k}-n^{2k}}{r-n^{2}}$
where $r=3\sum_{i=0}^{q-1}s_{i}n_{i}m_{i}p_{i}$.
Proof.
We can count the arithmetic cost by considering performing this restriction recursively: we perform $A$ arithmetic operations on matrices of size $n^{k-1}\times n^{k-1}$ at the first step, so the arithmetic operation count at this step is bounded by $An^{2(k-1)}$. At the second step we have at most $r$ lots of $n^{k-2}\times n^{k-2}$ matrices that we need to construct, and then use linear combinations to get outputs; this means the arithmetic operations at this recursive step is bounded by $Arn^{2(k-2)}$. Continuing this, we see that at the $j$-th recursive step, we have at most $Ar^{j-1}n^{2(k-j)}$ arithmetic operations. Summing all of these, we get that the arithmetic cost is bounded by the geometric sum
$An^{2(k-1)}+Arn^{2(k-2)}+\ldots+Ar^{k-1}=A\frac{r^{k}-n^{2k}}{r-n^{2}}.$
Combining these results we can show a complexity bound for Algorithm 1 using the restrictions as described in the proof of Theorem 2.
Given a restriction
$\langle n,n,n\rangle\leq_{\phi}\bigoplus_{i=0}^{q-1}s_{i}\odot(\langle n_{i},m_{i},p_{i}\rangle\oplus\langle m_{i},p_{i},n_{i}\rangle\oplus\langle p_{i},n_{i},m_{i}\rangle)$
we will show a bound on the number of operations to multiply $N\times N$ matrices using the restriction constructed from $\phi^{\otimes k}$ in the proof of Thereom 2. From the proof of Theorem 1 and Lemma 3 we can derive a bound on the leading coefficient $L_{k}=O(n^{4k})$ and from the discussion below Theorem 2 it follows that $\omega_{k}=\omega_{0}+O(1/k)$. Combined this gives that the algorithm takes at most $C_{1}n^{4k}N^{\omega_{0}+\frac{C_{2}}{k}}$ operations to multiply $N\times N$ matrices, for some positive constants $C_{1},C_{2}$. By taking $k=\frac{1}{2}\sqrt{\frac{C_{2}\log N}{\log n}}$, we can multiply $N\times N$ matrices in
$\operatorname{O}\left(N^{\omega_{0}}\exp\left(4\sqrt{C_{2}\log N\log n}\right)\right)$
operations.
Since we used several crude estimates, the bound we obtain here is likely far from tight. Therefore, we take the same approach as other literature (Karstadt and Schwartz, 2017; Schwartz et al., 2024; Schwartz and Zwecher, 2025) and use the leading coefficient computed assuming that every step of the recursion is performed exactly all the way down to $1\times 1\times 1$. Under this assumption, Algorithm 1 multiplies $N\times N$ matrices in at most
$\left(\frac{A(n+m+p)}{-3nmp+\sum_{i=0}^{q-1}s_{i}(n_{i}m_{i}p+n_{i}mp_{i}+nm_{i}p_{i})}+1\right)N^{\omega_{0}}+\operatorname{O}(N^{2})$
operations, though one should note that, unlike in standard Strassen-like algorithms, there is in general no input matrix for which this assumption holds. This is the leading coefficient and exponent we will be using to compare our algorithm to others. We expect that the $\exp\left(\operatorname{O}(\sqrt{\log N})\right)$ term could be dropped, but this requires a more careful analysis.
If we wanted to implement our algorithm then we would not go to a tensor power and instead apply Algorithm 1 directly using the original restriction and its cyclic permutations. As pointed out before, the computation performed in both cases is essentially the same. To guarantee that the recursive calls do not become too rectangular and thus require excessive zero padding, we can split rectangular subproblems into several square subproblems. Doing so every $k$ recursive steps, would achieve a similar complexity as the procedure described in the proof of Theorem 2. However, we are not restricted to this rigid procedure and could use a more refined heuristic to treat rectangular subproblems, like using different restrictions that have a rectangular matrix multiplication on the left side.
One should also note that we could use almost the same algorithm to make use of restrictions of the form $k\odot\langle n,n,n\rangle\leq\bigoplus_{i=0}^{q-1}s_{i}\odot\langle n_{i},m_{i},p_{i}\rangle$, which correspond to compute several disjoint matrix multiplications at once. So it gives an algorithmic version of the generalized ASI in the case where all matrix multiplication tensors on the left side have the same shape. Using the same approach, we find that the algorithm obtained from $k\odot\langle n,n,n\rangle\leq\bigoplus_{i=0}^{q-1}s_{i}\odot\langle n_{i},m_{i},p_{i}\rangle$ has a complexity of
$\left(\frac{A(n+m+p)}{-3nmpk+\sum_{i=0}^{q-1}s_{i}(n_{i}m_{i}p+n_{i}mp_{i}+nm_{i}p_{i})}+1\right)N^{\omega_{0}}+\operatorname{O}(N^{2})$
where $\omega_{0}$ is such that
$kn^{\omega_{0}}=\sum_{i=0}^{q-1}s_{i}(n_{i}m_{i}p_{i})^{\omega_{0}/3},$
again assuming that all recursive calls can be done exactly all the way down to $1\times 1\times 1$.
5. Specific Decompositions
We have already used the Moosbauer and Poole’s tensor decomposition of $\langle 6,6,6\rangle$ in the form $137\odot\langle 1\rangle\oplus 8\odot\langle 2,1,1\rangle$ as illustrating example. According to Theorem 2, this decomposition leads to an exponent $2.80496$, slightly better than the exponent $2.80754$ announced by Moosbauer and Poole and than Strassen’s exponent $2.80735$, though not as good as the exponent $2.7925$ of the decompositions of Novikov et al. (Novikov et al., 2025b) and Dumas, Pernet, and Sedoglavic (Dumas et al., 2025). We have repeated the computation of Moosbauer and Pool in order to generate further decompositions of $\langle 6,6,6\rangle$ in the hope to find some that contain more copies of $\langle 2,1,1\rangle$. The best we found contains $18$ such copies and thus leads to an exponent $2.8016$. The leading coefficient for this decomposition turns out to be $7.22$, compared to $6$ for Strassen’s algorithm.
We have also repeated the computation of Moosbauer and Pool for $\langle 5,5,5\rangle$ in the hope to find a decomposition with a better structure. The best we found leads to an exponent $2.809$, slightly higher than Strassen’s exponent.
In a next step, for various choices of $n,m,p$, we have searched for decompositions of $\langle n,m,p\rangle$ that contain copies of $\langle 1,2,2\rangle$ and/or $\langle k,1,1\rangle$ for $k=2,3,4$, or permutations of these. This search procedure is described in more detail below. The results are summarized in Table 1. In this table, $\omega_{\text{rank}}$ refers to the exponent reported in Sedoglavic’s table (Sedoglavic, 2025), $\omega_{\text{new}}$ refers to the exponent obtained via Theorem 2, $L$ is the leading coefficient given by (2) and $A$ the number of additions as determined by the software of Mårtensson and Wagner  (Mårtensson and Wagner, 2025), and the structure indicator $1^{u}2^{v}3^{w}4^{x}$ means that we have found a decomposition consisting of $u$ copies of $\langle 1\rangle$, altogether $v$ tensors of rank 2 (i.e., $\langle 1,1,2\rangle$, $\langle 1,2,1\rangle$, or $\langle 2,1,1\rangle$), altogether $w$ tensors of rank 3 (a permutation of $\langle 1,1,3\rangle$), and $x$ tensors of rank 4 (either a permutation of $\langle 1,1,4\rangle$ or a permutation of $\langle 1,2,2\rangle$).
$n$ $m$ $p$ $\omega_{\text{rank}}$ $\omega_{\text{new}}$ $L$ AA structure
4 4 4 2.7925 2.7925 11.7 341 $1^{48}$
2 2 2 2.8073 2.8073 6 15 $1^{7}$
6 6 6 2.8075 2.8016 7.22 691 $1^{117}\,2^{18}$
5 5 5 2.816 2.809 6.84 374 $1^{72}\,2^{6}\,3^{3}$
3 3 4 2.819 2.813 6.06 89 $1^{26}\,3$
2 2 8 2.884 2.815 6.48 91 $1\,3\,4^{6}$
2 2 6 2.874 2.816 5.88 55 $1\,2\,3^{6}$
2 3 8 2.859 2.818 8.45 182 $1^{2}\,2^{13}\,4^{3}$
2 4 4 2.820 2.818 6.50 88 $1^{24}\,2$
2 3 4 2.828 2.820 6.05 58 $1^{16}\,2^{2}$
2 4 6 2.839 2.823 6.59 142 $1^{19}\,2^{10}$
3 3 3 2.854 2.836 6.08 61 $1^{15}\,2^{2}\,4$
2 4 5 2.844 2.838 6.35 115 $1^{27}\,2^{3}$
2 2 4 2.856 2.838 5.60 33 $1^{8}\,2^{3}$
2 3 6 2.847 2.839 5.73 92 $1^{22}\,2^{4}$
2 2 7 2.898 2.840 5.90 70 $1\,2\,3^{2}\,4^{4}$
2 3 7 2.854 2.842 5.93 115 $1^{21}\,2^{7}$
2 2 5 2.894 2.850 5.28 39 $1^{2}\,2^{2}\,3^{4}$
2 2 3 2.895 2.883 5.48 24 $1^{7}\,2^{2}$
Table 1. New decompositions found for various matrix multiplication tensors $\langle n,m,p\rangle$. The first two rows are included for comparison only.
It turned out that no other case leads to an exponent smaller than Strassen’s, but we do see noticeable drops in the exponent compared to the previously best known exponent obtained from decompositions of the corresponding size. The decompositions claimed in the table are available electronically at
https://github.com/mkauers/matrix-multiplication/structured
In order to find these decompositions, we proceeded according to the following steps.
Step 1. In order to obtain a decomposition containing, say, $\langle 1,2,2\rangle$, we first apply a flip graph search (Kauers and Moosbauer, 2022, 2023a; Moosbauer and Poole, 2025; Kauers and Wood, 2025b; Perminov, 2025) to the tensor
$\sum_{i=1}^{n}\sum_{j=1}^{m}\sum_{k=1}^{p}a_{i,j}\otimes b_{j,k}\otimes c_{k,i}$
$-a_{1,1}\otimes b_{1,1}\otimes c_{1,1}-a_{2,1}\otimes b_{1,1}\otimes c_{1,2}$
$-a_{1,2}\otimes b_{2,1}\otimes c_{1,1}-a_{2,2}\otimes b_{2,1}\otimes c_{1,2}$
in order to find a decomposition with as low rank as possible. Adding
$a_{1,1}\otimes b_{1,1}\otimes c_{1,1}+a_{2,1}\otimes b_{1,1}\otimes c_{1,2}$
$+a_{1,2}\otimes b_{2,1}\otimes c_{1,1}+a_{2,2}\otimes b_{2,1}\otimes c_{1,2}$
to the result yields a decomposition of $\langle n,m,p\rangle$ which contains a copy of $\langle 1,2,2\rangle$.
Compared to the computation time invested into this first step, the computation time of all the subsequent steps is negligible.
Step 2. Flip graph searches have so far only been employed in order to find decompositions of low rank. The same technique can however also be used to optimize other features of a decomposition. On the decompositions obtained from step 1, we performed a flip graph search with the aim of maximizing the number of copies of $\langle 1,1,2\rangle$, $\langle 1,2,1\rangle$, $\langle 2,1,1\rangle$ contained in it. Note that these patterns are quite easy to detect. They amount to two components $a\otimes b\otimes c$ which have one factor in common. More generally, a copy of $\langle 1,1,k\rangle$ (or a permutation) corresponds to $k$ components sharing a common factor, so also these are easy to detect.
It must be noted however that in order to apply Algorithm 1, we must use a selection of copies of $\langle 1,1,k\rangle$, $\langle 1,\ell,1\rangle$, $\langle m,1,1\rangle$ that do not overlap. For example,
$a\otimes b\otimes c+a\otimes b^{\prime}\otimes c^{\prime}+a^{\prime\prime}\otimes b^{\prime}\otimes c^{\prime\prime}$
contains a copy of $\langle 2,1,1\rangle$ (because $a$ appears twice) as well as a copy of $\langle 1,2,1\rangle$ (because $b^{\prime}$ appears twice), but we must not use both of them because they overlap in $a\otimes b^{\prime}\otimes c^{\prime}$. The exponents in the structure indicator in Table 1 take this issue into account.
Step 3. To the decompositions obtained in Step 2, we next apply a number of random elements of de Groote’s symmetry group (de Groote, 1978a, b) in search for an orbit element with a small support. In principle, it would also be possible to minimize support with a flip graph search, but this approach would likely destroy the structure constructed in steps 1 and 2. In contrast, the structure is invariant under symmetries, and therefore preserved by the application of elements of the symmetry group.
Step 4. Up to this point, all computations are done over the field $\mathbb{Z}_{2}$. In the next step, we lift the coefficients of the decomposition to integers. There are various ways for doing this. Hensel lifting (Kauers and Moosbauer, 2023a; Gathen and Gerhard, 2003) was used in earlier papers about flip graph searches. However, as pointed out by Kemper (Kemper, 2025), the resulting coefficients tend to be more complicated than necessary. In several instances where Hensel lifting led to decompositions involving rational numbers with rather lengthy numerators and denominators, he was able to obtain a decomposition involving only the coefficients 1,0,1-1,0,1, starting from the same decomposition over $\mathbb{Z}_{2}$.
More critically, we must take care that the structure imposed on the decomposition during steps 1 and 2 is preserved during lifting. This is not automatically ensured. For example, the decomposition of $\langle 6,6,6\rangle$ presented by Moosbauer and Poole contains two copies of $\langle 1,1,1\rangle$ which become a copy of $\langle 1,1,2\rangle$ when coefficients are taken modulo $2$.
We continue to use Hensel lifting and address this issue by imposing additional constraints in order to ensure that the structure of the decomposition is preserved. In the present context, Hensel lifting leads to an underdetermined linear system over $\mathbb{Z}_{2}$, and the additional constraints can be easily encoded as additional linear equations which we append to this system. In the same way, we try to preserve the sparsity from Step 3 by imposing additional constraints so as to ensure that every zero in $\mathbb{Z}_{2}$ will remain a zero during the lifting. This may seem somewhat aggressive, but it worked surprisingly well and in many cases led to decompositions involving only the coefficients 1,0,1-1,0,1.
Step 5. Finally, we determine for each decomposition the number of additions needed to execute it. For this step, we employ software provided by Mårtensson and Wagner  (Mårtensson and Wagner, 2025).
6. Simulation
Since the bounds on the leading coefficient do not fully reflect the actual operation count of the algorithms, we simulate the recursive calls performed by our algorithm for different input sizes $N$ in order to get a more accurate idea of the runtime of the algorithm. Of course this approach does not take into account memory access costs and other practical considerations, but it does give a better estimate of the operation count than just the leading coefficient.
We compare Winograd’s variant of Strassen’s algorithm, which has a leading coefficient of $6$, to our algorithm using the decomposition of $\langle 6,6,6\rangle$ we found. The blue $\times$ and orange ++ in Figure 1 show the simulated operation counts for the algorithms, where for Moosbauer-Poole, we switch to Winograd-Strassen for $N<10^{4}$ and in both cases switch to the standard algorithm for $N<35$ as one would do in an actual implementation. The lines show the complexity estimate using the leading coefficient according to the formulas given above. The reason that our simulation shows lower operation counts than the reported bounds is that we switch to more efficient algorithms for small matrix sizes, while in the analysis we assumed that the recursion is performed exactly all the way down to $1\times 1$ matrices.
We can see that the results from our simulation do not form a smooth curve. This is due to the necessary zero padding, which makes the algorithms sensitive to the input size. In the simulations we start to see improvements over Strassen’s algorithm around matrix sizes of about $10^{6}$ and a consistent outperformance starting at $10^{10}$.
Figure 1. Simulated operation count for our algorithm and Strassen’s algorithm

### v2

@@
 I. Wood was supported by the Austrian FWF grant 10.55776/PAT8258123
-conference: ; ; $c$s: Computing methodologies Algebraic algorithms
+journalyear: The results in this paper are flawed. Please do not use them. We are working on a revised version. conference: ; ; $c$s: Computing methodologies Algebraic algorithms
 1. Introduction
@@
 Then we optimize the number of additions using DeGroote actions.
-These steps produce an algorithm that is valid only modulo $2$, thus we apply Hensel lifting to lift the algorithm to work over the integers, and finally we reduce the number of additions further by Mårtensson and Wagner’s common subexpression elimination method (Mårtensson and Wagner, 2025).
+These steps produce an algorithm that is valid only modulo 22, thus we apply Hensel lifting to lift the algorithm to work over the integers, and finally we reduce the number of additions further by Mårtensson and Wagner’s common subexpression elimination method (Mårtensson and Wagner, 2025).
 We find improvements in the exponent for several small base cases including $3\times 3$ matrix multiplication, where we reduce the exponent from Laderman’s $2.854$ (Laderman, 1976) to $2.836$.
@@
 First, it reduces the multiplication of two $n\times n$ matrices to multiplying two $2\times 2$ matrices whose entries are $\frac{n}{2}\times\frac{n}{2}$ matrices using block matrix multiplication.
-Then it computes this $2\times 2$ block matrix product using only $7$ multiplications of the subblocks by linearly combining them in a clever way.
+Then it computes this $2\times 2$ block matrix product using only 77 multiplications of the subblocks by linearly combining them in a clever way.
 Using this strategy recursively gives an algorithm to multiply $n\times n$ matrices using $\operatorname{O}(n^{\log_{2}(7)})$ operations in $R$.
@@
 $\Phi(a_{i_{2}j_{1}}\otimes b_{j_{2}k_{1}}\otimes c_{k_{2}i_{1}})(\mathbf{A},\mathbf{B})=\mathbf{A}_{i_{2}j_{1}}\mathbf{B}_{j_{2}k_{1}}\varGamma_{i_{1}k_{2}}$
-where $\mathbf{A}\in A,\mathbf{B}\in B$, $\mathbf{A}_{ij},\mathbf{B}_{ij}$ represent the $(i,j)$-th entry of $\mathbf{A}$ and $\mathbf{B}$ respectively and $\varGamma_{ij}\in C$ is the matrix with all zeros except for a $1$ at the $(i,j)$-th entry.
+where $\mathbf{A}\in A,\mathbf{B}\in B$, $\mathbf{A}_{ij},\mathbf{B}_{ij}$ represent the $(i,j)$-th entry of $\mathbf{A}$ and $\mathbf{B}$ respectively and $\varGamma_{ij}\in C$ is the matrix with all zeros except for a 11 at the $(i,j)$-th entry.
 Note the $c_{ij}$ corresponds to $\mathbf{C}_{ji}$ in the product matrix; this is a standard convention to keep the cyclic nature of matrix multiplication.
@@
 We will denote this number by $A(\phi)$ if the restriction is given by $\phi$.
-The homomorphism for a restriction is in general not unique, but in our case every restriction is witnessed by an explicit homomorphism, so we just write $A$ when the homomorphism is clear from the context.
+The homomorphism for a restriction is in general not unique, but in our case every restriction is witnessed by an explicit homomorphism, so we just write AA when the homomorphism is clear from the context.
 Using this definition, Strassen’s original algorithm has $A=18$.
@@
 for all $N\geq n$.
-We will show by induction on $N$ that
+We will show by induction on NN that
 $T(N)\leq LN^{\omega_{0}}-dN^{2}$
@@
 Usually the leading coefficient bounds are computed in the idealized case where every recursive call is assumed to be done exactly.
-In this case, the leading coefficient of Strassen’s algorithm is $7$, the general formula being $L\leq\frac{A}{r-n^{2}}+1$ (Karstadt and Schwartz, 2017).
+In this case, the leading coefficient of Strassen’s algorithm is 77, the general formula being $L\leq\frac{A}{r-n^{2}}+1$ (Karstadt and Schwartz, 2017).
 Theorem 3 is very important, however, all improvements since 1981 to the upper bound on $\omega$ have come from the following stronger theorem due to Schönhage.
@@
 $\kern-1.0pt\oplus 8^{2}\cdot 137\odot(\langle 1,2,2\rangle\oplus\langle 2,1,2\rangle\oplus\langle 2,2,1\rangle)\oplus 8^{3}\odot\langle 2,2,2\rangle.$
-While we have to compute $\langle 1,1,2\rangle,\langle 1,2,2\rangle$ and their cyclic permutations by the standard algorithm, using $2$ and $4$ multiplications respectively, we can compute $\langle 2,2,2\rangle$ using Strassen’s algorithm with $7$ multiplications.
+While we have to compute $\langle 1,1,2\rangle,\langle 1,2,2\rangle$ and their cyclic permutations by the standard algorithm, using 22 and 44 multiplications respectively, we can compute $\langle 2,2,2\rangle$ using Strassen’s algorithm with 77 multiplications.
 In the language of tensor restrictions we use $\langle 1,1,2\rangle\leq\langle 2\rangle$, $\langle 1,2,2\rangle\leq\langle 4\rangle$ and $\langle 2,2,2\rangle\leq\langle 7\rangle$, to conclude that $\langle 216,216,216\rangle\leq\langle 3581065\rangle$.
@@
 Since $n^{\prime}>c$, we can ensure that the denominator is positive, so $L$ is always well defined.
-We will prove by induction on $N$ that
+We will prove by induction on NN that
 (1) $T(N)\leq LN^{\omega_{0}}-LcN^{\omega_{0}-1}-dN^{2}$
@@
 Proof.
-We can count the arithmetic cost by considering performing this restriction recursively: we perform $A$ arithmetic operations on matrices of size $n^{k-1}\times n^{k-1}$ at the first step, so the arithmetic operation count at this step is bounded by $An^{2(k-1)}$.
+We can count the arithmetic cost by considering performing this restriction recursively: we perform AA arithmetic operations on matrices of size $n^{k-1}\times n^{k-1}$ at the first step, so the arithmetic operation count at this step is bounded by $An^{2(k-1)}$.
 At the second step we have at most $r$ lots of $n^{k-2}\times n^{k-2}$ matrices that we need to construct, and then use linear combinations to get outputs; this means the arithmetic operations at this recursive step is bounded by $Arn^{2(k-2)}$.
@@
 The best we found contains $18$ such copies and thus leads to an exponent $2.8016$.
-The leading coefficient for this decomposition turns out to be $7.22$, compared to $6$ for Strassen’s algorithm.
+The leading coefficient for this decomposition turns out to be $7.22$, compared to 66 for Strassen’s algorithm.
 We have also repeated the computation of Moosbauer and Pool for $\langle 5,5,5\rangle$ in the hope to find a decomposition with a better structure.
@@
 The results are summarized in Table 1.
-In this table, $\omega_{\text{rank}}$ refers to the exponent reported in Sedoglavic’s table (Sedoglavic, 2025), $\omega_{\text{new}}$ refers to the exponent obtained via Theorem 2, $L$ is the leading coefficient given by (2) and $A$ the number of additions as determined by the software of Mårtensson and Wagner  (Mårtensson and Wagner, 2025), and the structure indicator $1^{u}2^{v}3^{w}4^{x}$ means that we have found a decomposition consisting of $u$ copies of $\langle 1\rangle$, altogether $v$ tensors of rank 2 (i.e., $\langle 1,1,2\rangle$, $\langle 1,2,1\rangle$, or $\langle 2,1,1\rangle$), altogether $w$ tensors of rank 3 (a permutation of $\langle 1,1,3\rangle$), and $x$ tensors of rank 4 (either a permutation of $\langle 1,1,4\rangle$ or a permutation of $\langle 1,2,2\rangle$).
+In this table, $\omega_{\text{rank}}$ refers to the exponent reported in Sedoglavic’s table (Sedoglavic, 2025), $\omega_{\text{new}}$ refers to the exponent obtained via Theorem 2, $L$ is the leading coefficient given by (2) and AA the number of additions as determined by the software of Mårtensson and Wagner  (Mårtensson and Wagner, 2025), and the structure indicator $1^{u}2^{v}3^{w}4^{x}$ means that we have found a decomposition consisting of $u$ copies of $\langle 1\rangle$, altogether $v$ tensors of rank 2 (i.e., $\langle 1,1,2\rangle$, $\langle 1,2,1\rangle$, or $\langle 2,1,1\rangle$), altogether $w$ tensors of rank 3 (a permutation of $\langle 1,1,3\rangle$), and $x$ tensors of rank 4 (either a permutation of $\langle 1,1,4\rangle$ or a permutation of $\langle 1,2,2\rangle$).
 $n$ $m$ $p$ $\omega_{\text{rank}}$ $\omega_{\text{new}}$ $L$ AA structure
@@
 This is not automatically ensured.
-For example, the decomposition of $\langle 6,6,6\rangle$ presented by Moosbauer and Poole contains two copies of $\langle 1,1,1\rangle$ which become a copy of $\langle 1,1,2\rangle$ when coefficients are taken modulo $2$.
+For example, the decomposition of $\langle 6,6,6\rangle$ presented by Moosbauer and Poole contains two copies of $\langle 1,1,1\rangle$ which become a copy of $\langle 1,1,2\rangle$ when coefficients are taken modulo 22.
 We continue to use Hensel lifting and address this issue by imposing additional constraints in order to ensure that the structure of the decomposition is preserved.
@@
 6. Simulation
-Since the bounds on the leading coefficient do not fully reflect the actual operation count of the algorithms, we simulate the recursive calls performed by our algorithm for different input sizes $N$ in order to get a more accurate idea of the runtime of the algorithm.
+Since the bounds on the leading coefficient do not fully reflect the actual operation count of the algorithms, we simulate the recursive calls performed by our algorithm for different input sizes NN in order to get a more accurate idea of the runtime of the algorithm.
 Of course this approach does not take into account memory access costs and other practical considerations, but it does give a better estimate of the operation count than just the leading coefficient.
-We compare Winograd’s variant of Strassen’s algorithm, which has a leading coefficient of $6$, to our algorithm using the decomposition of $\langle 6,6,6\rangle$ we found.
+We compare Winograd’s variant of Strassen’s algorithm, which has a leading coefficient of 66, to our algorithm using the decomposition of $\langle 6,6,6\rangle$ we found.
 The blue $\times$ and orange ++ in Figure 1 show the simulated operation counts for the algorithms, where for Moosbauer-Poole, we switch to Winograd-Strassen for $N<10^{4}$ and in both cases switch to the standard algorithm for $N<35$ as one would do in an actual implementation.